OpenStudy (anonymous):

i need help: Express the simultaneous equations 2x+ky=7 and 4x-9y=4 in matrix form. Given that the above equations have no solutions/ find the possible value of k. i am also wondering what does it mean when it say "the equations have no solution" ?

OpenStudy (anonymous):

Pleas help me :/

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

You know matrix multiplication? So.... \[\huge A=\left[\begin{matrix}2 & k \\ 4 & -9\end{matrix}\right]\] And let's consider this matrix... \[\huge \left[\begin{matrix}x \\ y\end{matrix}\right]\]

terenzreignz (terenzreignz):

What do you get when you multiply these two? \[\huge \left[\begin{matrix}2 & k \\ 4 & -9\end{matrix}\right]\left[\begin{matrix}x \\ y\end{matrix}\right]\]

OpenStudy (anonymous):

i will get the same equation as above

terenzreignz (terenzreignz):

That's right :) So it all boils down to this... \[\huge \left[\begin{matrix}2 & k \\ 4 & -9\end{matrix}\right]\left[\begin{matrix}x \\ y\end{matrix}\right]=\left[\begin{matrix}7 \\ 4\end{matrix}\right]\] This seem about right?

OpenStudy (anonymous):

yup

terenzreignz (terenzreignz):

Okay, so we want this to have no solution, right? For convenience, let's set \[\huge z=\left[\begin{matrix}x \\ y\end{matrix}\right] \ \ \ \ ; \ \ \ c=\left[\begin{matrix}7 \\ 4\end{matrix}\right]\] ok?

OpenStudy (anonymous):

okayyy....

terenzreignz (terenzreignz):

So we can rewrite this as... \[\huge Az=c\] See how much simpler that looks? :P

OpenStudy (anonymous):

yeah..

terenzreignz (terenzreignz):

So.. you now have your matrix form... Az=c just replace the matrices as necessary with their actual values. Now we're interested in when this system has no solution... any idea how to do it? :)

OpenStudy (anonymous):

that is exactly what i don't know :s

terenzreignz (terenzreignz):

Okay... forget about those for a bit, and remember matrices... specifically, you know of inverse matrices?

OpenStudy (anonymous):

yes i know

terenzreignz (terenzreignz):

So, what if A has an inverse? Then... \[\huge Az=c\]\[\huge A^{-1}Az=A^{-1}c\]\[\huge z=A^{-1}c\] Catch me so far?

OpenStudy (anonymous):

yes totally!

terenzreignz (terenzreignz):

So... you realise that IF A has an inverse, this WILL have a solution, right? :D

OpenStudy (anonymous):

maybe .. :P what am i suppose to say here? :s

terenzreignz (terenzreignz):

What you're supposed to say... sort of... is that if we don't want this system to have a solution, then A had better not have an inverse :) So... yeah, what values of k do you need so that the matrix A has no inverse?

OpenStudy (anonymous):

i am sorry i didn't get anything :/

terenzreignz (terenzreignz):

Well, let's go back to the matrix A \[\huge A=\left[\begin{matrix}2 & k \\ 4 & -9\end{matrix}\right]\] You've heard of determinants?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

A neat trick at finding out whether a (square) matrix has an inverse... It has an inverse IF AND ONLY IF its determinant is not zero. So... what's the determinant of A?

OpenStudy (anonymous):

-18-4k

terenzreignz (terenzreignz):

LOL that's right! Now, so that the matrix A has no inverse, this determinant must be equal to 0 -18 - 4k = 0 Now solve for k

OpenStudy (anonymous):

ahh i get it now so by no solution we mean that A should not have an inverse and that only happens when determinant is equal to 0 right? i am getting -4 1/2 is that right cause in my book the answer is 4 1/2 ?

terenzreignz (terenzreignz):

I'm sure you're right. Perhaps you overlooked a negative-sign in your book?

OpenStudy (anonymous):

no i didn't and that was one of the reasons i posted the question here i thought i was doing wrong -.-

terenzreignz (terenzreignz):

\[\huge A^{-1}=\left[\begin{matrix}\frac14 & \frac18 \\ \frac19 & \frac{-1}{18}\end{matrix}\right]\]

terenzreignz (terenzreignz):

That's what happens if k=4 1/2 A would have an inverse, and this system would have a solution.

OpenStudy (shubhamsrg):

You have my support @terenzreignz , all your logic is correct, but the inverse of A is wrong! -_- Nevermind, it doesn't matter! ^_^

OpenStudy (shubhamsrg):

-4 1/2 will give equal slopes and not 4 1/2

terenzreignz (terenzreignz):

Inverting matrices was never my strong point... but I stand by this inverse... Unless you have a better idea, @shubhamsrg ? :D

OpenStudy (anonymous):

yeah i really didn't get how you got this inverse?

OpenStudy (shubhamsrg):

1/4 -1/9 -1/8 -1/18 that should be the inverse

terenzreignz (terenzreignz):

I don't think so... Try multiplying it with A

OpenStudy (shubhamsrg):

leme try.. hmm

OpenStudy (anonymous):

@terenzreignz is right :)

OpenStudy (shubhamsrg):

I see where I went wrong, sorry buddy.

terenzreignz (terenzreignz):

Outvoted, @shubhamsrg @Farheen28 I used an online calculator (read: I cheated) There are ways to calculate the inverse of matrices, one way in particular that works with 2x2's

OpenStudy (shubhamsrg):

hmmm.. :|

terenzreignz (terenzreignz):

I mean, works really well with 2x2 matrices :D That involves first getting the adjoint of the matrix... know what that is?

OpenStudy (anonymous):

anyways thanks a ton @terenzreignz you explained everything brilliantly ! :D i now know everything clearly! adjoint is determinant if i am not wrong :/

terenzreignz (terenzreignz):

adjoint? I am confused with its actual definition, but for 2x2 matrices \[\large \left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\] Its adjoint is \[\large \left[\begin{matrix}d & -b \\ -c & a\end{matrix}\right]\] and to get the inverse of the original 2x2 matrix, divide each element of the adjoint by the determinant. You see this can't be done if the determinant is zero :D

OpenStudy (anonymous):

ah you cleared one more thing! GENIUS :D

terenzreignz (terenzreignz):

Nope... just lucky :D And this concludes another episode of "Terence Teaches Matrices" tune in next time, see you around XD

OpenStudy (anonymous):

hhahahah i so will tune in again :D